OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (14, -2) and is perpendicular to 7x – 10y = 18 y=-7/10x+9/5 is where i am stuck at?

OpenStudy (agentjamesbond007):

So now that you got the y-intercept form, you know that to get a perpendicular side, you invert and change the side of the slope so the equation is now y=10/7x+9/5 now, you need point-slope for it, plug in the coordinate with the new slope (y-y1)=m(x-x1) (y+2)=10/7(x-14) distribute and it will give you the y-intercept form